OpenStudy (anonymous):

Analysis Question: Prove that if x [\in] (-1,1) and t is between 0 and x (so that t and x have the same sign and |t| [\leq] |x| < 1 then: [large \mid {x-t over t+1} \mid \leq |x|]

OpenStudy (anonymous):

gah... The following are my tex errors: \[\large x \in (-1,1)\] \[\large \left| t \right| \leq \left| x \right| < 1\] \[\large \left| {x-t \over t+1} \right| \leq |x|\]

OpenStudy (experimentx):

|x-t| <= |x| + |x||t| |x|+|t| <= |x| + |x||t| |t| <= |x|

OpenStudy (experimentx):

assume |x-t| <= |x| + |x||t|

OpenStudy (anonymous):

AH, that last part is what helped make this click. I had something like that but i think that is key to doing this correctly.